# Identification of Semiparametric Panel Multinomial

### v2

@@ 3.2 Computation
Recall that our identification strategy is based on the logical implication of the event $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0$, so for identification purposes we are only interested in $\mathbf{\mathbbm{1}}\left\{\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0\right\}$, i.e., whether the event $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0$ occurs, but not in the exact magnitude of $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$. However, when $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ is close to zero, the estimator $\hat{\gamma}_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ is relatively more likely to have the wrong sign, so that the plug-in estimator $\mathbf{\mathbbm{1}}\left\{\hat{\gamma}_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0\right\}$ may induce an error of the size $1$. Hence the smoothing by $G\left(\cdot\right)$ helps down-weight the observations when $\hat{\gamma}_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ is close to zero and shrinks the magnitude of possible errors.
On the other hand, when $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ is positive and large so that $\mathbf{\mathbbm{1}}\left\{\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0\right\}$ can be estimated well, we do not care much about the magnitude of $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$, which does not provide additional identifying information. By setting $G$ to be bounded from above, we dampen the effects of large $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ at the same time, so that the numerical maximization of $\hat{Q}$ is not too sensitive to potential large but redundant variations in $\hat{\gamma}_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$.
@@
In the optimization of $\hat{Q}\left(\beta\right)$ over $\beta\in\mathbb{\mathbb{S}}^{D-1}$, we work with a reparameterization of $\mathbb{\mathbb{S}}^{D-1}$ with $\left(D-1\right)$ angles in spherical coordinates [^7] The idea and the motivations for using the angle-space reparameterization were also found in Manski and Thompson (1986), who however used only one angle parameter.. Specifically, define the angle space $\Theta$ by
@@
and the transformation $\theta\longmapsto\beta\left(\theta\right)$ by standard spherical coordinate trasnformation, we now instead solves the optimization of $\hat{Q}\left(\beta\left(\theta\right)\right)$ over $\Theta$, which we further equip with its natural geodesic metric $\rho_{\Theta}\left(\theta,\tilde{\theta}\right):=\arccos\left(\beta\left(\theta\right)^{{}^{\prime}}\beta\left(\tilde{\theta}\right)\right)$, which is strongly equivalent [^8] Two metrics $d_{1}$ and $d_{2}$ defined on some non-empty set $X$ are strongly equivalent if and only if there exist positive constants $c_{1}$ and $c_{2}$ such that for every $x,y\in X$, we have $c_{1}d_{1}\left(x,y\right)\leq d_{2}\left(x,y\right)\leq c_{2}d_{1}\left(x,y\right).$ to the (imported) Euclidean distance $\left\|\beta\left(\theta\right)-\beta\left(\tilde{\theta}\right)\right\|$.
This reparameterization $\left(\Theta,\rho_{\Theta}\right)$ enables us to exploit the compactness and convexity of the parameter space $\Theta=\left[-\pi,\pi\right)\times\left[-\frac{\pi}{2},\frac{\pi}{2}\right]^{D-2}$, which takes the form of a hyper-rectangle. First, $\left(\Theta,\rho_{\Theta}\right)$ preserves all topological structure of the unit sphere, and particularly inherits the compactness of $\left(\mathbb{\mathbb{S}}^{D-1},\left\|\cdot\right\|\right)$, automatically satisfying the compactness condition usually imposed for extremum estimation and making it numerically feasible to initiate a grid on the whole parameter space. Second, while the unit sphere $\mathbb{\mathbb{S}}^{D-1}$ is not convex, the new parameter space $\Theta$ becomes convex algebraically, making it computationally easy to define bisection points in the parameter space. Third, it also preserves the geometric structures of the sphere, including for instance the obvious observation that $-\pi$ and $\pi$ in the first coordinate of $\Theta$ should be treated as exactly the same point, or more rigorously, $\rho_{\Theta}\left(\left(\pi-\epsilon,\theta_{2},...,\theta_{D-1}\right),\left(-\pi,\theta_{2},...,\theta_{D-1}\right)\right)\to 0$ as $\epsilon\to 0$. This seemingly trivial property is nevertheless important in defining and interpreting whether certain parameter estimates converge asymptotically or not.
@@
With the angle reparameterization, we seek to numerically compute a conservative rectangular enclosure of $\arg\min\hat{Q}\left(\theta\right)$, deploying a bisection-style grid-search algorithm that recursively shrinks and refines an adaptive grid to any pre-chosen precision (as defined by $\rho_{\Theta}$). Unlike gradient-based local optimization algorithms, our adaptive grid algorithm handles well the built-in discreteness in our sample criterion function, which has zero derivative almost everywhere, while maintains global initial coverage over the whole parameter space. While a brute-force global search algorithm is the safest choice if the dimension of product characteristics $D$ is relatively small, our adaptive-grid algorithm performs significantly faster. The essential structure of our algorithm is laid out as follows, with a corresponding illustration in Figure 1.
Figure 1: An Adaptive-Grid Algorithm
@@
Step 2: Compute $\hat{Q}\left(\theta\right)$ for each $\theta\in\Theta^{\left(1\right)}$, and select all points in $\Theta^{\left(1\right)}$ with a criterion value below the $\alpha$th-quantile in $\hat{Q}\left(\Theta^{\left(1\right)}\right):=\left\{\hat{Q}\left(\theta\right):\theta\in\Theta^{\left(1\right)}\right\}$ into
$\underline{\Theta}^{\left(1\right)}:=\left\{\theta\in\Theta^{\left(1\right)}:\ \hat{Q}\left(\theta\right)\leq\mathrm{quantile}_{\alpha}\left(\hat{Q}\left(\Theta^{\left(1\right)}\right)\right)\right\}.$
Step 3: Take the enclosing rectangle of $\underline{\Theta}^{\left(1\right)}$, by defining $\underline{\theta}_{d}^{\left(1\right)}:=\mathrm{min}^{*}\underline{\Theta}_{d}^{\left(1\right)}$ and $\overline{\theta}_{d}^{\left(1\right)}:=\mathrm{max}^{*}\underline{\Theta}_{d}^{\left(1\right)},$ where $\underline{\Theta}_{d}^{\left(1\right)}:=\left\{\theta_{d}:\theta\in\underline{\Theta}^{\left(1\right)}\right\}$ for each $d=1,...,D-1$ and the operator $\mathrm{min}^{*}$ and $\mathrm{max}^{*}$ have standard definitions of $\min$ and $\max$ except for the first dimension $d=1$. For the first dimension, it is necessary to account for the underlying spherical geometry and the periodicity of angles, i.e. $\theta_{1}+2\pi\equiv\theta_{1}$ and in particular $-\pi\equiv\pi$. This, however, is largely a programming nuisance: whenever $\underline{\Theta}_{1}^{\left(1\right)}\subsetneq\Theta_{1}^{\left(1\right)}$ crosses over at $-\pi$ and $\pi$, we can add $2\pi$ to every $\theta_{1}\in\underline{\Theta}_{1}^{\left(1\right)}$ and obtain lower and upper bounds of $\underline{\Theta}_{1}^{\left(1\right)}+2\pi$, as illustrated in Figure 1.
@@
Step 5: Reiterate until refinement stops (falls below a certain numerical precision).
Note that the above is simply a sketch of our algorithm. [^9] Our algorithm relies heavily on the compactness and convexity of the angle space $\Theta$. Compactness allows us to start with a global grid over the whole parameter space for initial evaluations of the sample criterion function. At each step of recursion, the convexity of $\Theta$ enables us to conveniently refine the grid by separately cutting each coordinate of $\overline{\underline{\Theta}}^{\left(m\right)}$ into smaller pieces through simple division. To be conservative, we add in buffers at each step of refinement, keep track of both outer and inner boundaries of the lower-quantile set $\underline{\Theta}^{\left(m\right)}$, and make sure that the minimizers of the criterion functions at all computed points are indeed enclosed by the set returned in the end. We find the current algorithm to be conservative and perform reasonably well in our simulations.
@@ 4 General Econometric Framework of Multi-Index Single-Crossing Conditions
Our key identification strategy, and consequently the associated estimation method, apply much more widely beyong panel multinomial choice models. We now introduce a general econometric framework defined by multi-index single-crossing (MISC) conditions, and show how our proposed methods can be exploited in a wide range of models nested under the MISC condition framework.
Formally, let $\left(y_{i},X_{i}\right)_{i=1}^{n}$ be a random sample of data with $X_{i}$ distributed on the support ${\cal X}\subseteq\mathbb{R}^{d_{x}}$ and $y_{i}$ distributed on ${\cal{\cal Y}}\subseteq\mathbb{R}^{d_{y}}$. Let $h_{0}:{\cal X}\to\mathbb{R}$ denote a functional of the conditional distribution of $y_{i}$ given $X_{i}$ that is directly identified from data. For each of $j=1,...,J\in\mathbb{N}$, let $\phi_{j}:{\cal X}\to\mathbb{R}^{d_{w}}$ be some known transformation of $X_{i}$ for $j=1,...,J$, and define $W_{ij}:=\phi_{j}\left(X_{i}\right)$ with $W_{i}:=\left(W_{i1},...,W_{iJ}\right)$. Let $\theta_{0j}\in\Theta_{j}\subseteq\mathbb{R}^{d_{\theta_{j}}}$ be an unknown finite-dimensional parameter and write $\theta_{0}:=\left(\theta_{01}^{{}^{\prime}},...,\theta_{0J}^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta:=\times_{j=1}^{J}\Theta_{j}$.
@@
We say that $\left(h_{0},\theta_{0}\right)$ satisfy the (weak) multi-index single-crossing condition if, for any realization $x\in{\cal X}$ and $w=\phi\left(x\right)$, we have
$w_{j}^{{}^{\prime}}\theta_{0j}\geq 0,\ \forall j=1,...,J\quad$ $\Rightarrow\quad h_{0}\left(x\right)\geq 0,$
$w_{j}^{{}^{\prime}}\theta_{0j}\leq 0,\ \forall j=1,...,J\quad$ $\Rightarrow\quad h_{0}\left(x\right)\leq 0.$ (15)
The condition is said to be strict if the inequalities on the right hand side of (15) are strict.
In words, if the $J$ parametric indexes $w_{1}^{{}^{\prime}}\theta_{01}$, $w_{2}^{{}^{\prime}}\theta_{02}$, … and $w_{J}^{{}^{\prime}}\theta_{0J}$ are all positive, then the functional $h_{0}$ must be positive; if the $J$ indexes are all zero, then $h_{0}$ must be zero; if the $J$ indexes are all negative, then $h_{0}$ must be negative. Essentially, the MISC condition provides a parsimonious way to semiparametrically model how the multiple economic factors jointly affect a certain statistic of the relevant economic outcome. The MISC condition basically requires that, if all the relevant factors reach certain threshold, then the outcome statistic must also reach certain threshold. Such kind of requirement is often very easy to obtain in an economic or econometric model: while multiple factors in an economic model may interact with each other in potentially complicated manners and there might be many configurations of the factors that lead to ambiguous theoretical predictions, there are also often simple configurations that we understand reasonably well. Hence, the MISC condition imposes relatively weak requirements on the economic or econometric model of a specific problem, and thus provides a relatively general framework for semiparametric econometric analysis, where most of the modeling ingredients can be left nonparametric beyond the parametric indexes that encode different economic factors in a certain economic problem.
Clearly, the panel multinomial choice model considered in previous sections falls under the MISC condition framework. Specifically, focusing on a pair of time periods $\left(t,s\right)$ and a particular product $j_{0}$ for illustration, define $\theta_{0j}:=\beta_{0}$, $h_{0}\left({\bf X}_{i}\right):=\gamma_{j_{0},ts}\left({\bf X}_{i}\right)$, $W_{ij_{0}}:=X_{ij_{0}t}-X_{ij_{0}s}$ and $W_{ij}=-\left(X_{ijt}-X_{ijs}\right)$ for $j\neq j_{0}$. Then the MISC conditions (15) are satisfied under model 1 and Assumptions 2-3.
@@
where $y_{i}$ denotes whether consumer $i$ purchases a certain, $X_{i1}$ denotes a vector of covariates that influence the consumer’s utility from a product, and $X_{i2}$ denotes a vector of covariates that influence the consumer’s awareness of the product (such as advertising). Here $J=2$, $X_{i}:=\left(X_{i1},X_{i2}\right)$, $W_{i1}:=X_{i1}$, and $W_{i2}:=X_{i2}$. Let the functional $h_{0}$ be defined by $h_{0}\left(x\right):=\mathbb{E}\left[\left.y_{i}\right|X_{i}=x\right]-\frac{1}{4}.$ Then, under the conditional median restrictions $\text{med}\left(\left.u_{i}\right|X_{i}\right)=\text{med}\left(\left.v_{i}\right|X_{i}\right)=0$ and the conditional independence restriction $\left.u_{i}\perp v_{i}\right|X_{i}$, it can be shown that
@@
again satisfying the MISC condition.
@@
and suppose that one component of $W_{i}$, say, $W_{i1}$ is endogenous. Suppose that there exist exogenous instrumental variables $Z_{i}$ and define $\xi_{i}:=W_{i1}-Z_{i}^{{}^{\prime}}\gamma_{0}$ as the residual from the reduced-form linear projection of $W_{i1}$ on $Z_{i}$. Assume that the endogeneity between $\epsilon_{i}$ and $W_{i1}$ is captured by the following control function
@@
where $\lambda$ is an unknown increasing function with location normalization $\lambda\left(0\right)=0$, and the sign parameter $\alpha_{0}\in\left\{-1,1\right\}$ controls the direction of the monotonicity. The above can be viewed as an adaption of the binary choice model that merges the conditional median restriction in Manski (1975) with the control function approach in Blundell and Powell (2004): here we only impose the control function restriction on the conditional median instead of the whole distribution as in Blundell and Powell (2004). Then, writing $\overline{Z}_{i}:=\left(W_{i1},Z_{i}\right)$ and $\overline{\gamma}_{0}:=\left(-\alpha_{0},\alpha_{0}\gamma_{0}\right)^{{}^{\prime}}$, we have
$W_{i}^{{}^{\prime}}\beta_{0}>0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}>0\quad\Rightarrow\quad\mathbb{E}\left[\left.Y_{i}\right|W_{i},Z_{i}\right]$ $>\frac{1}{2}.$
and its “$<$” counterpart, which can be viewed as a MISC condition with $K=2,$ $h_{0}\left(W_{i},Z_{i}\right):=\mathbb{E}\left[\left.Y_{i}-\frac{1}{2}\right|W_{i},Z_{i}\right]$, $\phi_{1}\left(W_{i},Z_{i}\right):=W_{i}$, $\phi_{2}\left(W_{i},Z_{i}\right):=\left(W_{i1},Z_{i}\right)$, $\theta_{0,1}:=\beta_{0}$, and $\theta_{0,2}:=\overline{\gamma}_{0}$.
@@
Consider the following dyadic network formation model studied in Gao, Li, and Xu (2023), which generalizes the one studied in Graham (2017):
$\mathbb{E}\left[\left.y_{ij}\right|X_{i},X_{j},A_{i},A_{j}\right]\ =$ $\psi\left(w\left(X_{i},X_{j}\right)^{{}^{\prime}}\theta_{0},A_{i},A_{j}\right).$
Here $y_{ij}$ is a binary outcome indicating whether individuals $i$ and $j$ are linked in an undirected network, $X_{i}$ and $X_{j}$ are the individuals’ observable covariates, $w\left(X_{i},X_{j}\right)$ is a known pairwise transformation of individual covariates (with the leading example being $w_{h}\left(X_{i},X_{j}\right):=\left|X_{i,h}-X_{j,h}\right|$ for each coordinate $h=1,...,d_{x}$), $A_{i}$ and $A_{j}$ are unobserved individual degree heterogeneity terms, and $\psi:\mathbb{R}^{3}\to\mathbb{R}$ is an unknown function assumed to be multivariate increasing in all its three arguments. Specifically, fixing a particular pair of individuals $\overline{i}$ and $\overline{j}$ and two generic realizations $\overline{x},\underline{x}$ of $X_{i}$, it can be shown that, with
@@
the weak MISC condition is satisfied (under quite mild additional conditions):
@@
The approach proposed in Example 2 above can also be adapted to the following censored monotone transformation model with endogeneity

### v3

@@ 3.3 Computation
We now explain how we implement the semiparametric two-step estimation procedure proposed above. Since the model’s criterion function is possibly non-convex, standard gradient-based optimizers are susceptible to converging to local minima. To address this, we employ a multi-stage adaptive-grid search algorithm that explores the parameter space more robustly and aims to locate the global minimizer of the objective function. The code for our computation algorithm is publicly available on GitHub. [^13] https://github.com/mingliecon/GL_PMC
@@
Recall that our identification strategy is based on the logical implication of the event $\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)>0$. Thus, for identification purposes we are only interested in $\mathbf{\mathbbm{1}}\left\{\gamma_{\tilde{{\cal J}},t,s}({\bf x}_{ts})>0\right\}$, i.e., whether the event $\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)>0$ occurs, but not in the exact magnitude of $\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)$. However, when $\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)$ is close to zero, the estimator $\hat{\gamma}_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)$ is relatively more likely to have the wrong sign, so that the plug-in estimator $\mathbf{\mathbbm{1}}\left\{\hat{\gamma}_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)>0\right\}$ may induce a large error of magnitude $1$. Hence, the smoothing by $G$ helps down-weight the observations when $\hat{\gamma}_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)$ is close to zero and shrinks the magnitude of possible errors.
On the other hand, when $\gamma_{\tilde{{\cal J}},t,s}({\bf x}_{ts})$ is positive and large so that $\mathbf{\mathbbm{1}}\left\{\gamma_{\tilde{{\cal J}},t,s}({\bf x}_{ts})>0\right\}$ can be estimated well, the magnitude of $\gamma_{\tilde{{\cal J}},t,s}({\bf x}_{ts})$ itself does not provide additional identifying information. By setting $G$ to be bounded from above, we dampen the influence of large values of $\gamma_{\tilde{{\cal J}},t,s}({\bf x}_{ts})$, so that the numerical minimization of $\hat{Q}$ is less sensitive to potentially large but redundant variations in $\hat{\gamma}_{\tilde{{\cal J}},t,s}({\bf x}_{ts})$.
@@
To minimize $\hat{Q}(\beta)$ over $\beta\in\mathbb{\mathbb{S}}^{D-1}$, we work with a reparameterization of $\mathbb{\mathbb{S}}^{D-1}$ with D1D-1 angles in spherical coordinates. [^14] The idea and the motivation for using the angle-space reparameterization can also be found in Manski and Thompson (1986), who however use only one angle parameter. Specifically, define the angle space $\Theta$ by
@@
[…]
This reparameterization $\left(\Theta,\rho_{\Theta}\right)$ enables us to exploit the compactness and convexity of the parameter space $\Theta=\left[-\pi,\pi\right)\times\left[-\frac{\pi}{2},\frac{\pi}{2}\right]^{D-2}$, which takes the form of a hyper-rectangle. First, $\left(\Theta,\rho_{\Theta}\right)$ preserves all topological structures of the unit sphere, and particularly inherits the compactness of $\left(\mathbb{\mathbb{S}}^{D-1},\left\|\cdot\right\|\right)$, automatically satisfying the compactness condition usually imposed for extremum estimation and making it numerically feasible to initiate a grid on the whole parameter space. Second, while the unit sphere $\mathbb{\mathbb{S}}^{D-1}$ is not convex, the new parameter space $\Theta$ becomes convex algebraically, making it computationally easy to define bisection points in the parameter space. Third, $\left(\Theta,\rho_{\Theta}\right)$ preserves the geometric structures of the sphere, including, for instance, the obvious observation that $\pi$ and $\pi$ in the first coordinate of $\Theta$ should be treated as exactly the same point, or more rigorously, $\rho_{\Theta}\left(\left(\pi-\epsilon,\theta_{2},\ldots,\theta_{D-1}\right),\left(-\pi,\theta_{2},\ldots,\theta_{D-1}\right)\right)\to 0$ as $\epsilon\to 0$. This seemingly trivial property is nevertheless important in defining and interpreting whether certain parameter estimates converge asymptotically or not.
@@
With the angle reparameterization, we seek to numerically compute a conservative rectangular enclosure of $\arg\min\hat{Q}\left(\theta\right)$, deploying a bisection-style grid-search algorithm that recursively shrinks and refines an adaptive grid to any pre-chosen precision (as defined by $\rho_{\Theta}$). Unlike gradient-based local optimization algorithms, our adaptive grid algorithm handles the built-in discreteness in our sample criterion function, whose derivative is zero almost everywhere, while still maintaining global coverage over the entire parameter space. While a brute-force global search algorithm is the safest choice when the dimension of the product characteristics $D$ is relatively small, our adaptive-grid algorithm runs significantly faster. The essential structure of our algorithm is laid out as follows.
@@
$\underline{\Theta}^{\left(1\right)}:=\left\{\theta\in\Theta^{\left(1\right)}:\ \hat{Q}\left(\theta\right)\leq\mathrm{quantile}_{\alpha}\left(\hat{Q}\left(\Theta^{\left(1\right)}\right)\right)\right\}.$ (18)
Step 3: Take the enclosing rectangle of $\underline{\Theta}^{\left(1\right)}$, by defining $\underline{\theta}_{d}^{\left(1\right)}:=\mathrm{min}^{*}\underline{\Theta}_{d}^{\left(1\right)}$ and $\overline{\theta}_{d}^{\left(1\right)}:=\mathrm{max}^{*}\underline{\Theta}_{d}^{\left(1\right)},$ where $\underline{\Theta}_{d}^{\left(1\right)}:=\left\{\theta_{d}:\theta\in\underline{\Theta}^{\left(1\right)}\right\}$ for each $d=1,\ldots,D-1$ and the operator $\mathrm{min}^{*}$ and $\mathrm{max}^{*}$ have standard definitions of $\min$ and $\max$ except for the first dimension $d=1$. For the first dimension, it is necessary to account for the underlying spherical geometry and the periodicity of angles, i.e. $\theta_{1}+2\pi\equiv\theta_{1}$ and in particular $\pi\equiv\pi$. This, however, is largely a programming nuisance: whenever $\underline{\Theta}_{1}^{\left(1\right)}\subsetneq\Theta_{1}^{\left(1\right)}$ crosses over at $\pi$ and $\pi$, we can add $2\pi$ to every $\theta_{1}\in\underline{\Theta}_{1}^{\left(1\right)}$ and obtain lower and upper bounds of $\underline{\Theta}_{1}^{\left(1\right)}+2\pi$, as illustrated in Figure 1.
@@
Step 5: Iterate until refinement stops (falls below a certain numerical precision).
Note that the above is simply a sketch of our algorithm: see Appendix D and the documentation on GitHub for more implementation details. [^16] Our algorithm relies heavily on the compactness and convexity of the angle space $\Theta$. Compactness allows us to start with a global grid over the whole parameter space for initial evaluations of the sample criterion function. At each step of recursion, the convexity of $\Theta$ enables us to conveniently refine the grid by separately cutting each coordinate of $\overline{\underline{\Theta}}^{\left(m\right)}$ into smaller pieces through simple division. To be conservative, we add in buffers at each step of refinement, keep track of both outer and inner boundaries of the lower-quantile set $\underline{\Theta}^{\left(m\right)}$, and make sure that the minimizers of the criterion functions at all computed points are indeed enclosed by the set returned in the end. We find the current algorithm to be conservative and to perform well in our simulations.
This multi-stage approach is designed to balance computational feasibility with a robust search. The initial coarse search efficiently discards large, suboptimal regions of the parameter space, while the subsequent refinement and boundary identification stages provide a high-precision estimate in the most promising area. However, the algorithm’s performance is inherently tied to the selection of tuning parameters, particularly the initial grid size M_Step and the quantile used for pruning, which must be chosen carefully to ensure the global minimum is not discarded prematurely. Furthermore, it can get computationally intensive when the dimension of $\beta$ is high. We find that running 1,000 simulations for $D=3$ usually takes a few hours on modern computers, while for $D=4$ it may take up to a day.
@@ 4 General Econometric Framework of Multi-Index Single-Crossing Conditions
Our key identification strategy, and consequently the associated estimation method, apply more widely beyond panel multinomial choice models. We now introduce a general econometric framework defined by multi-index single-crossing (MISC) conditions, and show how our proposed methods can be exploited in a wide range of models nested under the MISC condition framework.
Formally, let $\left(y_{i},X_{i}\right)_{i=1}^{n}$ be a random sample of data with $X_{i}$ distributed on the support ${\cal X}\subseteq\mathbb{R}^{d_{x}}$ and $y_{i}$ distributed on ${\cal{\cal Y}}\subseteq\mathbb{R}^{d_{y}}$. Let $h_{0}:{\cal X}\to\mathbb{R}$ denote a functional of the conditional distribution of $y_{i}$ given $X_{i}$ that is directly identified from data. For each of $j=1,\ldots,J\in\mathbb{N}$, let $\phi_{j}:{\cal X}\to\mathbb{R}^{d_{\theta_{j}}}$ be some known transformation of $X_{i}$, and define $W_{ij}:=\phi_{j}\left(X_{i}\right)$ with $W_{i}:=\left(W_{i1},\ldots,W_{iJ}\right)$. Let $\theta_{0j}\in\Theta_{j}\subseteq\mathbb{R}^{d_{\theta_{j}}}$ be an unknown finite-dimensional parameter and write $\theta_{0}:=\left(\theta_{01}^{{}^{\prime}},\ldots,\theta_{0J}^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta:=\times_{j=1}^{J}\Theta_{j}$.
@@
We say that $\left(h_{0},\theta_{0}\right)$ satisfy the (weak) multi-index single-crossing condition if, for any realization $x\in{\cal X}$ and $w=\phi\left(x\right)$,
$w_{j}^{{}^{\prime}}\theta_{0j}\geq 0,\ \forall j=1,\ldots,J\quad$ $\Rightarrow\quad h_{0}\left(x\right)\geq 0,$
$w_{j}^{{}^{\prime}}\theta_{0j}\leq 0,\ \forall j=1,\ldots,J\quad$ $\Rightarrow\quad h_{0}\left(x\right)\leq 0.$ (19)
The condition is said to be strict if the inequalities on the right-hand side of (19) are strict.
In words, the MISC condition states that if all the $J$ parametric indices $w_{1}^{{}^{\prime}}\theta_{01}$, $w_{2}^{{}^{\prime}}\theta_{02}$, $\ldots$, and $w_{J}^{{}^{\prime}}\theta_{0J}$ are (weakly) positive, then the functional $h_{0}$ must be (weakly) positive; if the $J$ indices are all zero, then $h_{0}$ must be zero; if the $J$ indices are all negative, then $h_{0}$ must be negative. Essentially, the MISC condition provides a parsimonious way to semiparametrically model how the multiple economic factors jointly affect a certain statistic of the relevant economic outcome. The MISC condition basically requires that, if all the relevant factors reach certain thresholds, then the outcome statistics must also reach certain thresholds. Such requirements are often easy to obtain in an economic or econometric model: while multiple factors in an economic model may interact with each other in potentially complicated manners and there might be many configurations of the factors that lead to ambiguous theoretical predictions, there are also often simple configurations that we understand reasonably well. Hence, the MISC condition imposes only mild requirements on the underlying economic or econometric model for the problem and thus provides a general framework for semiparametric econometric analysis, in which most of the modeling ingredients can be left nonparametric except for the parametric indices that capture different economic factors in the problem.
Clearly, the panel multinomial choice model considered in previous sections falls under the MISC condition framework. Specifically, focusing on a pair of time periods $\left(t,s\right)$ and a particular product $j_{0}$ for illustration, define $\theta_{0j}:=\beta_{0}$, $h_{0}\left({\bf X}_{i}\right):=\gamma_{j_{0},ts}\left({\bf X}_{i}\right)$, $W_{ij_{0}}:=X_{ij_{0}t}-X_{ij_{0}s}$ and $W_{ij}:=-\left(X_{ijt}-X_{ijs}\right)$ for $j\neq j_{0}$. Then, the MISC condition (19) is satisfied under model (1) and Assumptions 13.
@@
where $y_{i}$ denotes whether consumer $i$ purchases a certain product or not, $X_{i1}$ denotes a vector of covariates that influences the consumer’s utility from a product, and $X_{i2}$ denotes a vector of covariates that affects the consumer’s awareness of the product (e.g., advertising). Here, we have $J=2$, $X_{i}:=\left(X_{i1},X_{i2}\right)$, $W_{i1}:=X_{i1}$, and $W_{i2}:=X_{i2}$. Define $h_{0}\left(x\right):=\mathbb{E}\left[\left.y_{i}\right|X_{i}=x\right]-\frac{1}{4}.$ Then, under the conditional median restrictions $\text{med}\left(\left.u_{i}\right|X_{i}\right)=\text{med}\left(\left.v_{i}\right|X_{i}\right)=0$ and the conditional independence restriction $\left.u_{i}\perp v_{i}\right|X_{i}$, it is true that
@@
satisfying the MISC condition.
@@
and let one component of $W_{i}$, say, $W_{i1}$ be endogenous. Suppose that there exists a vector of instrumental variables $Z_{i}$ and define $\xi_{i}:=W_{i1}-Z_{i}^{{}^{\prime}}\gamma_{0}$ as the residual from the reduced-form linear projection of $W_{i1}$ on $Z_{i}$. Assume that the endogeneity between $\epsilon_{i}$ and $W_{i1}$ is captured by the following control function
@@
where $\lambda$ is an unknown increasing function with location normalization $\lambda\left(0\right)=0$, and the sign parameter $\alpha_{0}\in\left\{-1,1\right\}$ controls the direction of the monotonicity. The above can be viewed as an adaptation of the binary choice model that combines the conditional median restriction in Manski (1975) with the control function approach in Blundell and Powell (2004): here we only impose the control function restriction on the conditional median instead of the whole distribution as in Blundell and Powell (2004). Then, writing $\overline{Z}_{i}:=\left(W_{i1},Z_{i}\right)$ and $\overline{\gamma}_{0}:=\left(-\alpha_{0},\alpha_{0}\gamma_{0}\right)^{{}^{\prime}}$, we have
$W_{i}^{{}^{\prime}}\beta_{0}>0,\ \overline{Z}_{i}^{{}^{\prime}}\overline{\gamma}_{0}>0\quad\Rightarrow\quad\mathbb{E}\left[\left.Y_{i}\right|W_{i},Z_{i}\right]$ $>\frac{1}{2}$
and its “<<” counterpart, which can be viewed as a MISC condition with $K=2,$ $h_{0}\left(W_{i},Z_{i}\right):=\mathbb{E}\left[\left.Y_{i}-\frac{1}{2}\right|W_{i},Z_{i}\right]$, $\phi_{1}\left(W_{i},Z_{i}\right):=W_{i}$, $\phi_{2}\left(W_{i},Z_{i}\right):=\left(W_{i1},Z_{i}\right)$, $\theta_{0,1}:=\beta_{0}$, and $\theta_{0,2}:=\overline{\gamma}_{0}$.
@@
Consider the dyadic network formation model of Gao, Li, and Xu (2023), which extends Graham (2017) to a semiparametric setting:
$\mathbb{E}\left[\left.y_{ij}\right|X_{i},X_{j},A_{i},A_{j}\right]\ =\$$\psi\left(w\left(X_{i},X_{j}\right)^{{}^{\prime}}\theta_{0},A_{i},A_{j}\right).$
Here $y_{ij}$ is a binary outcome indicating whether individuals $i$ and $j$ are linked in an undirected network, $X_{i}$ and $X_{j}$ are the individuals’ observable covariates, $w(X_{i},X_{j})$ is a known pairwise transformation of individual covariates (with the leading example being $w_{h}\left(X_{i},X_{j}\right):=\left|X_{i,h}-X_{j,h}\right|$ for each coordinate $h=1,\ldots,d_{x}$), $A_{i}$ and $A_{j}$ are unobserved individual degree heterogeneity terms, and $\psi:\mathbb{R}^{3}\to\mathbb{R}$ is an unknown function assumed to be increasing in all its three arguments. Specifically, fixing a particular pair of individuals $(\overline{i},\overline{j})$ and two realizations $\overline{x},\underline{x}$ of $X_{i}$, it can be shown that, with
@@
the weak MISC condition is satisfied under mild conditions:
@@
The approach proposed in Example 2 above can also be adapted to the following censored monotone transformation model with endogeneity: